# Capturing reduced-order quantum many-body dynamics out of equilibrium via neural ordinary differential equations

### v1

@@ 3.2 Stability of the prediction
In the previous section, we focused on the short- to medium-term forecast. In Fig. 5, we consider the long-term prediction. The prediction diverges after a prediction time of about $t_{\text{pred}}=25-30\,J^{-1}$. Interestingly, the same behavior is observed in the analytical time-dependent two-particle reduced density matrix theory [donsa2023Nonequilibrium]. To stabilize the time-dependent two-particle reduced density matrix method, several necessary $N$-representability conditions are enforced during integration by means of a purification procedure [donsa2023Nonequilibrium, garrod1964Reduction, coleman1963Structure, mazziotti2012Structure, pescoller_projective_2025]. Specifically, the positive semi-definiteness of the two-particle reduced density matrix and the two-hole reduced density matrix are imposed. The matrix elements of the 2HRDM $Q_{12}(t)$ are given by $Q^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}=\langle\Psi(t)|a_{j_{1}\uparrow}a_{j_{2}\downarrow}a_{i_{1}\uparrow}^{\dagger}a_{i_{2}\downarrow}^{\dagger}|\Psi(t)\rangle$. Empirically, it has been found that this purification procedure stabilizes the time-dependent two-particle reduced density matrix very well in all parameter regions, where the time-local reconstruction functionals produce a sufficiently small error in the overall propagation [pescoller_projective_2025].
Figure 5: The occupation number $n_{1}$ of site 1 of the six-site Fermi-Hubbard model described in Eq. (1), with $V=1.0$ and $U=3.1$, is visualized over time by the blue solid line. The occupation number is calculated from the two-particle reduced density matrix. The predicted occupation number of four hyper-parameter optimized models trained on different losses are shown as well, all trained on the first 3000 time steps ($30\,J^{-1}$) of the time series. The orange dashed line was trained on the MSE loss only, the light gray dash-dotted line additionally on the constraint loss for the trace $\mathcal{L}_{\text{tr}}$, and the dark gray dotted line on the MSE and the negative eigenvalue loss $\mathcal{L}_{\text{psd}}$. The prediction starts from $D_{12}(t=40\,J^{-1})$, hence this is an extrapolation.
The question arises whether we can enforce these conditions on the neural ordinary differential equation as well? Do they stabilize the model’s long-term predictions? How do they perform compared to the unconstrained model? To enforce constraints in machine learning, there are, in principle, two options: i) weak constraints enforce the conditions via the loss function during training, while ii) hard constraints are built into the architecture of the model itself, preventing the model from violating these constraints. In the following, we will focus on weak constraints and their results.
@@ 3.2.1 Weak constraints
Due to their positive semi-definiteness, the negative eigenvalues $\lambda_{j}^{t}$ for both the two-particle reduced density matrix $D_{12}(t)$ and the two-hole reduced density matrix $Q_{12}(t)$ at time $t$ are penalized via
@@
The rectified linear unit (ReLU) is squared to obtain a continuous derivative. Thus, we have two additional terms in the loss, one for the two-particle reduced density matrix and one for the two-hole reduced density matrix. Similary, the trace conservation of the two-particle reduced density matrix over time
@@
for the two-particle reduced density matrix $D_{12}(t)$ and analogously for the two-hole reduced density matrix $Q_{12}(t)$.
For the following analysis, we focus on the parameter configuration V=1.0JV=1.0\,J and U=3.1JU=3.1\,J because of its dynamics, which is neither too slow nor too fast. Additionally, it is in the correlated regime, where the prediction should, in principle, work, but the performance in Fig. 3(d) is not perfect yet ($C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}=0.6899$). First, we optimize the model’s hyperparameters, resulting in much better performance: $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}=0.9194$. This supports the previous statement that the fluctuations in the performance in Fig. 3(c) and (d) could be reduced by hyper-parameter optimization. Next, we train three models on different combinations of the constraints, one on the trace constraint only, one on the eigenvalue constraint only, and one on both. Again, we optimize the hyperparameters of each model for a fair comparison. As depicted in Fig. 5, these constraints do not stabilize the long-term prediction. In fact, they all start to diverge at a similar prediction length, as defined in Eq. (17), $25\,J^{-1}<t_{\text{pred}}<30\,J^{-1}$. In Fig. 6 the Pearson correlation coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ and the different loss terms over the length of the prediction are visualized.
Figure 6: The Pearson correlation coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ (a), the MSE loss (b), the eigenvalue loss for the two-particle reduced density matrix $\mathcal{L}_{\text{psd,2RDM}}$ (c), and the two-hole reduced density matrix $\mathcal{L}_{\text{psd,2HRDM}}$ (d), as well as the trace loss for the two-particle reduced density matrix $\mathcal{L}_{\text{tr,2RDM}}$ (e) and the two-hole reduced density matrix $\mathcal{L}_{\text{tr,2HRDM}}$ (f) for models trained on different losses are all visualized over the length of the prediction. The blue solid line represents the unconstrained model, which was trained only on MSE loss. Models that had either trace constraint or negative eigenvalue constraint as additional loss terms are depicted as an orange dashed line and a light gray dashed-dotted line, respectively. The dark gray dotted line indicates the model trained with all of the previously mentioned loss terms.
The same divergent behavior is observed in the Pearson correlation. All models show similar performance across the various loss metrics, except for trace loss, where models trained with this constraint perform about two orders of magnitude better. However, this only leads to a minimally longer stable forecast length in the case of the trace constraint model, while there is no improvement for the model trained with both constraints. In general, the losses for the two-particle reduced density matrix are lower than those for the two-hole reduced density matrix, which is to be expected since the model is only trained directly with the two-particle reduced density matrix data, while the two-hole reduced density matrix is derived from it. Interestingly, for short-term predictions up to a prediction length of $t_{\text{pred}}=1700\,\text{steps}=17\,J^{-1}$ the unconstrained model performs best, but only slightly. This can be seen in the loss plots of the MSE and the eigenvalue constraint.
@@ 4 Discussion and Conclusion
Non-equilibrium dynamics in the Fermi-Hubbard model is Markovian on the level of the full many-body Schrödinger equation. Intricate non-Markovian behavior arises upon reduction of degrees of freedom, e.g. see Ref. [Coppola2025Learning]. Tracing out degrees of freedom from the full wave function leads to reduced density matrices such as the two-particle reduced density matrix. In this work, we showed how the neural ordinary differential equation serves as a model‑agnostic diagnostic tool to identify the degree of Markovianity of the two-particle reduced density matrix $D_{12}(t)$ dynamics. Its ability to reproduce the exact trajectories without any explicit knowledge about the three-particle reduced density matrix confirms that, in regimes where the Pearson coefficient $C_{\Delta_{12},\Delta_{123,\text{K}}}$ is large, the dynamics is effectively governed by a time‑local functional of the instantaneous two‑particle cumulant $\Delta_{12}(t)$. This means that no memory is required for the propagation. In the anti‑correlated regime, where $C_{\Delta_{12},\Delta_{123,\text{K}}}$ becomes negative, the neural ordinary differential equation fails to predict the evolution, indicating that no simple time‑local reconstruction of the three‑particle cumulant $\Delta_{123}(t)$ exists. In other words, memory is required.
@@
An analysis of the buildup of genuine three‑particle correlations $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}$ reveals that the magnitude of $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}$ is the primary factor limiting the applicability of current reconstruction functionals. Regions with $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}<0.65$ coincide with accurate neural ordinary differential equation forecasts and low reconstruction errors, whereas larger values correlate with the breakdown of both the neural ordinary differential equation and the time-dependent two-particle reduced density matrix method.
@@
These observations suggest a natural extension of the neural ordinary differential equation framework: by endowing the vector field with explicit history dependence – e.g., through finite‑time convolutions – one can empirically determine the minimal memory length required for an accurate reconstruction of $\Delta_{123}(t)$. Such a memory‑aware neural ordinary differential equation would provide quantitative guidance for developing non‑local reconstruction functionals that incorporate the necessary temporal kernels.
@@
During training, we imposed weak physical constraints (norm conservation and positive semi-definiteness) via additional loss terms. Although the norm constraint improves predictions slightly, the constraints did not yield a noticeable enhancement of long‑time stability. Enforcing hard constraints is more challenging for neural ordinary differential equations because they learn derivatives, thus symmetries that apply on the level of the integrated observable cannot be implemented straightforwardly into the architecture of the model. Recent proposals for constrained neural differential equations as in Ref. [white2023Stabilized] offer promising routes to incorporate such hard constraints.

### v2

@@ 3.2 Stability of the prediction
In the previous section, we focused on the short- to medium-term forecast. In Fig. 6, we consider the long-term prediction. The prediction diverges after a prediction time of about $t_{\text{pred}}=25-30\,J^{-1}$. We note that Van de Walle et al. [98], who report state-of-the-art extrapolation capabilities, demonstrate predictions only up to $t_{\text{pred}}=8\,J^{-1}$, with considerably fewer oscillations than those encountered in Fig. 6. Interestingly, the same behavior is observed in the analytical TD2RDM theory [34]. To stabilize the TD2RDM method, several necessary $N$-representability conditions are enforced, if broken, during integration by means of a purification procedure [34, 38, 29, 63, 77]. Specifically, the positive semi-definiteness of the 2RDM and the two-hole reduced density matrix (2HRDM) are imposed. The matrix elements of the 2HRDM $Q_{12}(t)$ are given by $Q^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}=\langle\Psi(t)|a_{j_{1}\uparrow}a_{j_{2}\downarrow}a_{i_{1}\uparrow}^{\dagger}a_{i_{2}\downarrow}^{\dagger}|\Psi(t)\rangle$. Empirically, it has been found that this purification procedure stabilizes the TD2RDM very well in all parameter regions, where the time-local reconstruction functionals produce a sufficiently small error in the overall propagation [77].
Figure 6: The occupation number $n_{1}$ of site 1 of the six-site Fermi-Hubbard model described in Eq. (1), with $V=1.0$ and $U=3.1$, is visualized over time by the blue solid line. The occupation number is calculated from the 2RDM. The predicted occupation number of four hyper-parameter optimized models trained on different losses are shown as well, all trained on the first 3000 time steps ($30\,J^{-1}$) of the time series. The orange dashed line was trained on the MSE loss only, the light gray dash-dotted line additionally on the constraint loss for the trace $\mathcal{L}_{\text{tr}}$, and the dark gray dotted line on the MSE and the negative eigenvalue loss $\mathcal{L}_{\text{psd}}$. The prediction starts from $D_{12}(t=40\,J^{-1})$, hence this is an extrapolation.
The question arises whether we can enforce these conditions on the neural ODE as well? Do they stabilize the model’s long-term predictions? How do they perform compared to the unconstrained model? To enforce constraints in ML, there are, in principle, two options: i) weak constraints enforce the conditions via the loss function during training, while ii) hard constraints are built into the architecture of the model itself, preventing the model from violating these constraints. In the following, we will focus on weak constraints and their results.
@@ 3.2.1 Weak constraints
Due to their positive semi-definiteness, the negative eigenvalues $\lambda_{j}^{t}$ for both the 2RDM $D_{12}(t)$ and the 2HRDM $Q_{12}(t)$ at time $t$ are penalized via
@@
The rectified linear unit (ReLU) is squared to obtain a continuous derivative.
Figure 7: The Pearson correlation coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ (a), the MSE loss (b), the eigenvalue loss for the 2RDM $\mathcal{L}_{\text{psd,2RDM}}$ (c), and the 2HRDM $\mathcal{L}_{\text{psd,2HRDM}}$ (d), as well as the trace loss for the 2RDM $\mathcal{L}_{\text{tr,2RDM}}$ (e) and the 2HRDM $\mathcal{L}_{\text{tr,2HRDM}}$ (f) for models trained on different losses are all visualized over the length of the prediction. The blue solid line represents the unconstrained model, which was trained only on MSE loss. Models that had either trace constraint or negative eigenvalue constraint as additional loss terms are depicted as an orange dashed line and a light gray dashed-dotted line, respectively. The dark gray dotted line indicates the model trained with all of the previously mentioned loss terms.
Thus, we have two additional terms in the loss, one for the 2RDM and one for the 2HRDM. Similarly, the trace conservation of the 2RDM over time
@@
for the 2RDM $D_{12}(t)$ and analogously for the 2HRDM $Q_{12}(t)$.
For the following analysis, we focus on the parameter configuration V=1.0JV=1.0\,J and U=3.1JU=3.1\,J because of its dynamics, which is neither too slow nor too fast. Additionally, it is in the correlated regime, where the prediction should, in principle, work, but the performance in Fig. 3(d) is not optimized ($C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}=0.6899$). First, we optimize the model’s hyperparameters, resulting in much better performance: $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}=0.9194$. This supports the previous statement that the fluctuations in the performance in Fig. 3(c) and (d) could be reduced by hyper-parameter optimization. Next, we train three models on different combinations of the constraints, one on the trace constraint only, one on the eigenvalue constraint only, and one on both. Again, we optimize the hyperparameters of each model for a fair comparison. As depicted in Fig. 6, these constraints do not stabilize the long-term prediction. In fact, they all start to diverge at a similar prediction length, as defined in Eq. (18), $25\,J^{-1}<t_{\text{pred}}<30\,J^{-1}$. In Fig. 7 the Pearson correlation coefficient $C_{D_{12}^{\text{pred}},D_{12}^{\text{target}}}$ and the different loss terms over the length of the prediction are visualized. The same divergent behavior is observed in the Pearson correlation. All models show similar performance across the various loss metrics, except for trace loss, where models trained with this constraint perform about two orders of magnitude better. However, this only leads to a minimally longer stable forecast length in the case of the trace constraint model, while there is no improvement in the Pearson correlation coefficient (see Fig. 7(a)) for the model trained with both constraints. In general, the losses for the 2RDM are lower than those for the 2HRDM, which is to be expected since the model is only trained directly with the 2RDM data, while the 2HRDM is derived from it. Interestingly, for short-term predictions up to a prediction length of $t_{\text{pred}}=1700\,\text{steps}=17\,J^{-1}$ the unconstrained model performs best, but only slightly. This can be seen in the loss plots of the MSE and the eigenvalue constraint.
@@ 4 Discussion and Conclusion
Non-equilibrium dynamics in the Fermi-Hubbard model is Markovian on the level of the full many-body Schrödinger equation. Intricate non-Markovian behavior arises upon reduction of degrees of freedom, e.g. see Ref. [30]. Tracing out degrees of freedom from the full wave function leads to reduced density matrices such as the 2RDM. In this work, we showed how the neural ODE serves as a model‑agnostic diagnostic tool to identify the degree of Markovianity of the 2RDM $D_{12}(t)$ dynamics. Its ability to reproduce the exact trajectories without any explicit knowledge about the 3RDM confirms that, in regimes where the Pearson coefficient between the two- and three-particle cumulants $C_{\Delta_{12},\Delta_{123,\text{K}}}$ is large, the dynamics is effectively governed by a time‑local functional of the instantaneous two‑particle cumulant $\Delta_{12}(t)$. This means that no memory is required for the propagation. In the anti‑correlated regime, where $C_{\Delta_{12},\Delta_{123,\text{K}}}$ becomes negative, the neural ODE fails to predict the evolution, indicating that no simple time‑local reconstruction of the three‑particle cumulant $\Delta_{123}(t)$ exists. In other words, memory is required.
We want to emphasize that the neural ODE should not replace exact diagonalization or other physical algorithms, but rather tests whether the dynamics can be treated as Markovian. Since Markovianity is a widely used assumption in many methods and theories across different fields, we argue that a dedicated tool to verify this assumption is valuable, even at the computational cost observed here. At the same time, the proposed framework is not restricted to quantum systems. It can be applied to other types of time-series data, including settings with substantially lower dimensionality, where smaller models are sufficient and computational cost can be reduced significantly. Therefore, if exact reference data are available, we can ask a concrete and practically relevant question before committing to a given approximation scheme: is this specific assumption (Markovianity) justified for the dynamics at hand? As discussed in the introduction, existing Markovianity tests are either restricted to specific settings (e.g., one-dimensional reaction coordinates [8], open-system quantum channels [103], or stationary stochastic processes [21, 111, 89]) or provide only a binary accept/reject verdict. The neural ODE framework fills this gap by offering a general, assumption-free diagnostic that yields a continuous measure of effective Markovianity and is directly applicable to any temporal coherent time-series data, including the high-dimensional, deterministic, non-stationary dynamics of reduced density matrices.
Improving the quality of the extrapolated dynamics via neural ODEs benefits remarkably strongly from exploring the hyper-parameter space to find the near-optimal parameters. Hyper-parameter optimization works even better to obtain robust results than model-ensemble averaging. We achieve an extrapolation of several thousand time-steps, corresponding to a prediction time between $25\,J^{-1}<t_{\text{pred}}<30\,J^{-1}$ in our model.
@@
An analysis of the buildup of genuine three‑particle correlations $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}$ reveals that the magnitude of $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}$ is the primary factor limiting the applicability of current reconstruction functionals. Regions with $\overline{\delta\Delta^{\uparrow\uparrow\downarrow}_{123,\text{K}}}<0.65$ coincide with accurate neural ODE forecasts and low reconstruction errors, whereas larger values correlate with the breakdown of both the neural ODE and the TD2RDM method.
@@
These observations suggest a natural extension of the neural ODE framework: by endowing the vector field with explicit history dependence – e.g., with an explicit history channel, where the vector field depends on both the current state and a learned memory state, $\dot{y}(t)=f_{\theta}\big(y(t),m(t)\big)$ with $\dot{m}(t)=g_{\phi}\big(m(t),y(t)\big)$, or equivalently a finite-history formulation $\dot{y}(t)=f_{\theta}\big(y(t),y(t-\tau_{1}),\ldots,y(t-\tau_{k})\big)$ – one can empirically determine the minimal memory length required for an accurate reconstruction of $\Delta_{123}(t)$. Such a memory‑aware neural ODE would provide quantitative guidance for developing non‑local reconstruction functionals that incorporate the necessary temporal kernels.
@@
During training, we imposed weak physical constraints (norm conservation and positive semi-definiteness) via additional loss terms. Although the norm constraint improves predictions slightly, the constraints did not yield a noticeable enhancement of long‑time stability. Enforcing hard constraints is more challenging for neural ODEs because they learn derivatives, thus symmetries that apply on the level of the integrated observable cannot be implemented straightforwardly into the architecture of the model. Recent proposals for constrained neural differential equations as in Ref. [102] offer promising routes to incorporate such hard constraints.
@@ Acknowledgments
Data availability
The data to reproduce the experiments described in the article are publicly available on DaRUS [37].